OpenStudy (anonymous):

Question: The third and seventh terms of a geometric sequence are 3/4 and 12 respectively. A. Find the 10th term B. What term is equal to 3072 Answers: A. ±96 B. 15th Please tell me how we arrive at those answers. I'm a bit confused because the common ratio I get is quite weird (it's in fraction) and my calculator can't even solve it. I first got the difference between the 3rd and 12th terms and got 45/4 and then divided that by 4 to get the common difference of 45/16. Tell me what's wrong. thank you :)

OpenStudy (texaschic101):

Look at this and scroll down to #8. I have trouble find the first term if it is not given... let me try.... an = a1 * r^(n-1) 12 = 3/4 * r^(5-1) 12 = 3/4 * r^4 48/3 = r^4 16 = r^4 2 = r does that look right to you ?

OpenStudy (texaschic101):

oops....I forgot this... http://www.regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm

OpenStudy (anonymous):

So did you regard the 3rd term as 1st and solve it?

OpenStudy (texaschic101):

it is not solved, I just got the difference as being 2

OpenStudy (anonymous):

Yes, but the 2 there is a common ratio right?

OpenStudy (texaschic101):

yes....thats what I got....I can solve for the 10th term, if you think the common ratio is correct. Like I said, I am not good at finding the first term if it is not given, so I am not sure if the common ratio is correct.I think so, but not 100% sure

OpenStudy (anonymous):

So regarding 3rd term as 1st, in order to find out the 100th term, do we do 3/4(2^98-1) So 3/4(2^97)? Shouldn't that be the answer to A IF the common ratio is correct?

OpenStudy (anonymous):

I solved B, I'm a bit confused with A though. I think that's right but I don't know how to simplify that 3/4(2^97)

OpenStudy (texaschic101):

wait....to find the first term you kinda work backwards....3rd term is 3/4, common ratio is 2.....3/4 * 1/2 = 3/8 <-- 2nd term...3/8 * 1/2 = 3/16 <-- first term.. let me check... 3/16 x 2 = 6/16 = 3/8 (2nd term 3/8 x 2 = 6/8 = 3/4 (3rd term 6/8 x 2 = 12/8 = 6/4 = 3/2 (4th term 3/2 x 2 = 6/2 = 3 (5th term 3 x 2 = 6 (6th term 6 x 2 = 12 (7th term yes.....2 is common difference and 1st term is 3/16

OpenStudy (texaschic101):

an = a1 * r^(n - 1) n = 10 a1 = 3/16 r = 2 a10 = 3/16 ( 2^(10 - 1) a10 = 3/16 * 2^9 a10 = 3/16 * 512 a10 = 1536/16 a10 = 96 you might better get a second opinion on this

OpenStudy (anonymous):

haha alright, thank you so much :)

OpenStudy (texaschic101):

I tried....I really hope it is correct :)